### v2

# DiReCT: Diagnostic Reasoning for Clinical Notes via Large Language Models

Abstract
Large language models (LLMs) have recently showcased remarkable capabilities, spanning a wide range of tasks and applications, including those in the medical domain. Models like GPT-4 excel in medical question answering but may face challenges in the lack of interpretability when handling complex tasks in real clinical settings. We thus introduce the diagnostic reasoning dataset for clinical notes (DiReCT), aiming at evaluating the reasoning ability and interpretability of LLMs compared to human doctors. It contains 511 clinical notes, each meticulously annotated by physicians, detailing the diagnostic reasoning process from observations in a clinical note to the final diagnosis. Additionally, a diagnostic knowledge graph is provided to offer essential knowledge for reasoning, which may not be covered in the training data of existing LLMs. Evaluations of leading LLMs on DiReCT bring out a significant gap between their reasoning ability and that of human doctors, highlighting the critical need for models that can reason effectively in real-world clinical scenarios Code are available https://github.com/wbw520/DiReCT. Data will be released through PhysioNet..
22footnotetext: Corresponding author.
1 Introduction
Recent advancements of large language models (LLMs) [Zhao et al., 2023] have ushered in new possibilities and challenges for a wide range of natural language processing (NLP) tasks [Min et al., 2023]. In the medical domain, these models have demonstrated remarkable prowess [Anil et al., 2023, Han et al., 2023], particularly in medical question answering (QA) [Jin et al., 2021]. Leading-edge models, such as GPT-4 [OpenAI, 2023a], exhibit profound proficiency in understanding and generating text [Bubeck et al., 2023], even achieved high scores on the United States Medical Licensing Examination (USMLE) questions [Nori et al., 2023].
Despite the advancements, interpretability is critical, particularly in medical NLP tasks [Liévin et al., 2024]. Some studies assess this capability over medical QA [Pal et al., 2022, Li et al., 2023, Chen et al., 2024] or natural language inference (NLI) [Jullien et al., 2023]. Putting more attention on interpretability, they use relatively simple tasks as testbeds, taking short text as input. However, tasks in real clinical settings can be more complex [Gao et al., 2023a]. As shown in Figure 1, a typical diagnosis requires comprehending and combining various information, such as health records, physical examinations, and laboratory tests, for further reasoning of possible diseases in a step-by-step manner following the established guidelines. This observation suggests that both perception, or reading, (e.g., finding necessary information in medical record) and reasoning (determining the disease based on the observations) should be counted when evaluating interpretability in LLM-based medical NLP tasks.
For a more comprehensive evaluation of LLMs for supporting diagnosis in a more realistic setting, we propose a Diagnostic Reasoning dataset for Clinical noTes (DiReCT). The task basically is predicting the diagnosis from a clinical note of a patient, which is a collection of various medical records, written in natural language. Our dataset contains 511 clinical notes spanning 25 disease categories, sampled from a publicly available database, MIMIC-IV [Johnson et al., 2023]. Each clinical note undergoes fine-grained annotation by professional physicians. The annotators (i.e., the physicians) are responsible for identifying the text, or the observation, in the note that leads to a certain diagnosis, as well as the explanation. The dataset also provides a diagnostic knowledge graph based on existing diagnostic guidelines to facilitate more consistent annotations and to supply a model with essential knowledge for reasoning that might not be encompassed in its training data.
To underscore the challenge offered by our dataset, we evaluate a simple AI-agent based baseline [Xi et al., 2023, Tang et al., 2023] that utilizes the knowledge graph to decompose the diagnosis into a sequence of diagnoses from a smaller number of observations. Our experimental findings indicate that current state-of-the-art LLMs still fall short of aligning well with human doctors.
Contribution. DiReCT offers a new challenge in diagnosis from a complex clinical note with explicit knowledge of established guidelines. This challenge aligns with a realistic medical scenario that doctors are experiencing. In the application aspect, the dataset facilitates the development of a model to support doctors in diagnosis, which is error-prone [Middleton et al., 2013, Liu et al., 2022]. From the technical aspect, the dataset can benchmark models’ ability to read long text and find necessary observations for multi-evidence entailment tree reasoning. As shown in Figure 3, this is not trivial because of the variations in writing; superficial matching does not help, and medical knowledge is vital. Meanwhile, reasoning itself is facilitated by the knowledge graph. The model does not necessarily have the knowledge of diagnostic guidelines. With this choice, the knowledge graph explains the reasoning process, which is also beneficial when deploying such a diagnosis assistant system in practical uses.
Figure 1: When a patient is admitted, an initial consultation takes place to collect subjective information. Subsequent observations may then require further examination to confirm the diagnosis.
2 Related Works
Natural language explanation. Recent advancements in NLP have led to significant achievements [Min et al., 2023]. However, existing models often lack explainability, posing potential risks [Danilevsky et al., 2020, Gurrapu et al., 2023]. Numerous efforts have been made to address this challenge. One effective approach is to provide a human-understandable plain text explanation alongside the model’s output [Camburu et al., 2018, Rajani et al., 2019]. Another strategy involves identifying evidence within the input that serves as a rationale for the model’s decisions, aligning with human reasoning [DeYoung et al., 2020]. Expanding on this concept, [Jhamtani and Clark, 2020] introduces chain-structured explanations, given that a diagnosis can demand multi-hop reasoning. This idea is further refined by ProofWriter [Tafjord et al., 2021] through a proof stage for explanations, and by [Zhao et al., 2021] through retrieval from a corpus. [Dalvi et al., 2021] proposes the entailment tree, offering more detailed explanations and facilitating inspection of the model’s reasoning. More recently, [Zhang et al., 2024] employed cumulative reasoning to tap into the potential of LLMs to provide explanation via a directed acyclic graph. Although substantial progress has been made, interpreting NLP tasks in medical domains remains an ongoing challenge [Liévin et al., 2024].
Benchmarks of interpretability in the medical domain Several datasets are designed to assess a model’s reasoning together with its interpretability in medical NLP (Table 1). MedMCQA [Pal et al., 2022] and other medical QA datasets [Li et al., 2023, Chen et al., 2024] provide plain text as explanations for QA tasks. NLI4CT [Jullien et al., 2023] uses clinical trial reports, focusing on NLI supported by multi-hop reasoning. N2N2 [Gao et al., 2022] proposes a summarization (Sum) task for a diagnosis based on multiple pieces of evidence in the input clinical note. NEJM CPC [Zack et al., 2023] interprets clinicians’ diagnostic reasoning as plain text for reasoning clinical diagnosis (CD). DR.BENCH [Gao et al., 2023b] aggregates publicly available datasets to assess the diagnostic reasoning of LLMs. Utilizing an multi-evidence entailment tree explanation, DiReCT introduces a more rigorous task to assess whether LLMs can align with doctors’ reasoning in real clinical settings.
Table 1: Comparison of existing datasets for medical reasoning tasks and ours. “t” and “w” mean tokens and words for the length of input, respectively.
Dataset Task Data Source Length Explanation # Cases
MedMCQA [Pal et al., 2022] QA Examination 9.93 t Plain Text 194,000
ExplainCPE [Li et al., 2023] QA Examination 37.79 w Plain Text 7,000
JAMA Challenge [Chen et al., 2024] QA Clinical Cases 371 w Plain Text 1,524
Medbullets [Chen et al., 2024] QA Online Questions 163 w Plain Text 308
N2N2 [Gao et al., 2022] Sum Clinical Notes 785.46 t Evidences 768
NLI4CT [Jullien et al., 2023] NLI Clinical Trail Reports 10-35 t Multi-hop 2,400
NEJM CPC [Zack et al., 2023] CD Clinical Cases - Plain Text 2,525
DiReCT (Ours) CD Clinical Notes 1074.6 t Entailment Tree 511
3 A benchmark for Clinical Notes Diagnosis
This section first detail clinical notes (Section 3.1). We also describes the knowledge graph that encodes existing guidelines (Section 3.2). Our task definition, which tasks a clinical note and the knowledge graph as input is given in Section 3.4. We then present our annotation process for clinical notes (Section 3.3) and the evaluation metrics (Section 3.5).
3.1 Clinical Notes
Clinical notes used in DiReCT are stored in the SOAP format [Weed, 1970]. A clinical note comprises four components: In the subjective section, the physician records the patient’s chief complaint, the history of present illness, and other subjective experiences reported by the patient. The objective section contains structural data obtained through examinations (inspection, auscultation, etc.) and other measurable means. The assessment section involves the physician’s analysis and evaluation of the patient’s condition. This may include a summary of current status, etc. Finally, the plan section outlines the physician’s proposed treatment and management plan. This may include prescribed medications, recommended therapies, and further investigations. A clinical note also includes a primary discharge diagnosis (PDD) in the assessment section.
DiReCT’s clinical notes are sourced from the MIMIC-IV dataset [Johnson et al., 2023] (PhysioNet Credentialed Health Data License 1.5.0), which encompasses over 40,000 patients admitted to the intensive care units. Each note contains clinical data for a patient. To construct DiReCT, we curated a subset of 511 notes whose PDDs fell within one of 25 disease categories $i$ in 5 medical domains.
In our task, a note $R=\{r\}$ is an excerpt of 6 clinical data in the subjective and objective sections (i.e., $|R|=6$): chief complaint, history of present illness, past medical history, family history, physical exam, and pertinent results. [^1] We excluded data, such as review system and social history, because they are often missing in the original clinical notes and are less relevant to the diagnosis. We also identified the PDD $d^{\star}$ associated with $R$ [^2] All clinical notes in DiReCT are related to only one PDD, and there is no secondary discharge diagnosis. The set of $d^{\star}$’s for all $R$’s collectively forms $\mathcal{D}^{\star}$. We manually removed any descriptions that disclose the PDD in $R$.
3.2 Diagnostic Knowledge Graph
Existing knowledge graphs for the medical domain, e.g., UMLS KG [Bodenreider, 2004], lack the ability to provide specific clinical decision support (e.g., diagnostic threshold, context-specific data, dosage information, etc.), which are critical for accurate diagnosis.
Our knowledge graphs $\mathcal{K}=\{\mathcal{K}_{i}\}_{i}$ is a collection of graph $\mathcal{K}_{i}$ for disease category $i$. $\mathcal{K}_{i}$ is based on the diagnosis criteria in existing guidelines (refer to supplementary material for details). $\mathcal{K}_{i}$’s nodes are either premise $p\in\mathcal{P}_{i}$ (medical statement, e.g., Headache is a symptom of) and diagnoses $d\in\mathcal{D}_{i}$ (e.g., Suspected Stroke). $\mathcal{K}_{i}$ consists of two different types of edges. One is premise-to-diagnosis edges $\mathcal{S}_{i}=\{(p,d)\}$, where $p\in\mathcal{P}_{i}$ and $d\in\mathcal{D}_{i}$; an edge is from $p$ to $d$. This edge represents the necessary premise $p$ to make a diagnosis $d$. We refer to them as supporting edges. The other is diagnosis-to-diagnosis edges $\mathcal{F}_{i}=\{(d,d^{\prime})\}$, where $d,d^{\prime}\in\mathcal{D}_{i}$ and the edge is from $d$ to $d^{\prime}$, which represents the diagnostic flow. These edges are referred to as procedural edges.
A disease category is defined according to an existing guideline, which starts from a certain diagnosis; therefore, a procedural graph $\mathcal{G}_{i}=(\mathcal{D}_{i},\mathcal{F}_{i})$ has only one root node and arbitrarily branches toward multiple leaf nodes that represent PDDs (i.e., the clinical notes in DiReCT are chosen to cover all leaf nodes of $\mathcal{G}_{i}$). Thus, $\mathcal{G}_{i}$ is a tree. We denote the set of the leaf nodes (or PDDs) as $\mathcal{D}^{\star}_{i}\subset\mathcal{D}_{i}$. The knowledge graph is denoted by $\mathcal{K}_{i}=(\mathcal{D}_{i},\mathcal{P}_{i},\mathcal{S}_{i},\mathcal{F}_{i})$.
Figure 2: A part of $\mathcal{K}_{i}$ for $i$ being Acute Coronary Syndromes.
Figure 2 shows a part of $\mathcal{K}_{i}$, where $i$ is Acute Coronary Syndromes (ACS). Premises in $\mathcal{P}_{i}$ and diagnoses in $\mathcal{D}_{i}$ are given in the blue and gray boxes, while PDDs in $\mathcal{D}^{\star}_{i}$ are ones without outgoing edges (i.e., STEMI-ACS and NSTEMI-ACS, and UA). The black and red arrows are edges in $\mathcal{S}$ and $\mathcal{F}$, respectively, where the black arrows indicate the supporting edges.
$\mathcal{K}$ serves two essential functions: (1) They serve as the gold standard for annotation, guiding doctors in the precise and uniform interpretation of clinical notes. (2) Our task also allows a model to use them to ensure the output from an LLM can be closely aligned with the reasoning processes of medical professionals.
3.3 Data Annotation
Let $d^{\star}\in\mathcal{D}^{\star}_{i}$ denote the PDD of disease category $i$ associated with $R$. We can find a subgraph $\mathcal{K}_{i}(d^{\star})$ of $\mathcal{K}_{i}$ that contains all ancestors of $d^{\star}$, including premises in $\mathcal{P}_{i}$. We also denote the set of supporting edges in $\mathcal{K}_{i}(d^{\star})$ as $\mathcal{S}_{i}(d^{\star})$. Our annotation process is, for each supporting edge $(p,d)\in\mathcal{S}_{i}(d^{\star})$, to extract observation $o\in\mathcal{O}$ in $R$ (highlighted text in the clinical note in Figure 3) and provide rationalization $z$ of this deduction why $o$ is a support for $d$ or corresponds to $p$ [^3] All annotations strictly follow the procedural flow in $\mathcal{K}_{i}$, and each observation is only related to one diagnostic node. If $R$ does not provide sufficient observations for the PDD (which may happen when a certain test is omitted), the annotators were asked to add plausible observations to $R$. This choice compromises the fidelity of our dataset to the original clinical notes, but we chose it for the completeness of the dataset. They form the explanation =$\mathcal{E}=\{(o,z,d)\}$ for $(R,d^{\star})$. This annotation process was carried out by 9 clinical physicians and subsequently verified for accuracy and completeness by three senior medical experts.
Table 2: Statistics of DiReCT.
Medical domain # cat. # samples $|\mathcal{D}_{i}|$ $|\mathcal{D}^{\star}_{i}|$ $|\mathcal{O}|$ Length
Cardiology 7 184 27 16 8.7 1156.6 t
Gastroenterology 4 103 11 7 4.3 1026.0 t
Neurology 5 77 17 11 11.9 1186.3 t
Pulmonology 5 92 26 17 10.7 940.7 t
Endocrinology 4 55 20 14 6.9 1063.5 t
Overall 25 511 101 65 8.5 1074.6 t
Table 2 summarizes statistics of our dataset. The second and third columns (“# cats.” and “# samples”) show the numbers of disease categories and samples in the respective medical domains. $|\mathcal{D}_{i}|$ and $|\mathcal{D}_{i}^{\star}|$ are the total numbers of diagnoses (diseases) and PDDs, summed over all diagnostic categories in the medical domain, respectively. $|\mathcal{O}|$ is the average number of annotated observations. “Length” is the average number of tokens in $R$.
Figure 3: An annotation sample of Heart Failure (HF). The left part is the clinical note alongside extracted observations by a doctor. The middle part outlines the steps of the rationale for the premise corresponding to each diagnostic node shown in the right part.
3.4 Task Definition
We propose two tasks with different levels of supplied external knowledge. The first task is, given $R$ and $\mathcal{G}$, to predict the associated PDD $d^{\star}$ and generate an explanation $\mathcal{E}$ that explains the model’s diagnostic procedure from $R$ to $d^{\star}$, i.e., letting $M$ denote a model:
$\hat{d}^{\star},\hat{\mathcal{E}}=M(R,\mathcal{G}),$ (1)
where $\hat{d}^{\star}\in\cup_{i}\mathcal{D}^{\star}_{i}$ and $\hat{\mathcal{E}}$ are predictions for the PDD and explanation, respectively. With this task, the knowledge of specific diagnostic procedures in existing guidelines can be used for prediction, facilitating interpretability. The second task takes $\mathcal{K}$ as input instead of $\mathcal{G}$, i.e.,:
$\hat{d}^{\star},\hat{\mathcal{E}}=M(R,\mathcal{K}).$ (2)
This task allows for the use of broader knowledge of premises for prediction. One may also try a task without any external knowledge.
3.5 Evaluation Metrics
We designed three metrics to quantify the predictive performance over our benchmark.
(1) Accuracy of diagnosis $\textit{Acc}^{\text{diag}}$ evaluates if a model can correctly identify the diagnosis. $\textit{Acc}^{\text{diag}}=1$ if $d^{\star}=\hat{d}$, and $\textit{Acc}^{\text{diag}}=0$ otherwise. The average is reported.
(2) Completeness of observations $\textit{Obs}^{\text{comp}}$ evaluates whether a model extracts all and only necessary observations for the prediction. Let $\mathcal{O}$ and $\hat{\mathcal{O}}$ denote the sets of observations in $\mathcal{E}$ and $\hat{\mathcal{E}}$, respectively. The metric is defined as $\textit{Obs}^{\text{comp}}=|\mathcal{O}\cap\hat{\mathcal{O}}|/|\mathcal{O}\cup\hat{\mathcal{O}}|$, where the numerator is the number of observations that are common in both $\mathcal{O}$ and $\hat{\mathcal{O}}$. [^4] We find the common observations with an LLM (refer to the supplementary material for more detail). This metric simultaneously evaluates the correctness of each observation and the coverage. To supplement it, we also report the precision $\textit{Obs}^{\text{pre}}$ and recall $\textit{Obs}^{\text{rec}}$, given by $\textit{Obs}^{\text{pre}}=|\mathcal{O}\cap\hat{\mathcal{O}}|/|\hat{\mathcal{O}}|$ and $\textit{Obs}^{\text{rec}}=|\mathcal{O}\cap\hat{\mathcal{O}}|/|\mathcal{O}|$.
(3) Faithfulness of explanations Faith evaluates if the diagnostic flow toward the PDD is fully supported by observations with faithful rationalizations. This involves establishing a one-to-one correspondence between deductions in the prediction and the ground truth. We use the correspondences established for computing $\textit{Obs}^{\text{comp}}$. Let $o\in\mathcal{O}$ and $\hat{o}\in\hat{\mathcal{O}}$ denote corresponding observations. This correspondence is considered successful if $z$ and $\hat{z}$ as well as $d$ and $\hat{d}$ associated with $o$ and $\hat{o}$ matches. Let $m(\mathcal{E},\hat{\mathcal{E}})$ denote the number of successful matches. We use the ratio of $m(\mathcal{E},\hat{\mathcal{E}})$ to $|\mathcal{O}\cap\hat{\mathcal{O}}|$ and $|\mathcal{O}\cup\hat{\mathcal{O}}|$ as evaluation metrics $\textit{Exp}^{\text{com}}$ and $\textit{Exp}^{\text{all}}$, respectively, to see failures come from observations or explanations and diagnosis.
4 Baseline
Figure 4: Pipeline of our baseline. The dotted line in the right-most boxes means deductions from an observation to a diagnosis.
Figure 4 shows an overview of our baseline with three LLM-based modules narrowing-down, perception, and reasoning (refer to the supplementary material for more details). The narrowing-down module $U$ takes $R$ as input to make a prediction $\hat{i}$ of the disease category, i.e., $\hat{i}=U(R)$.
Let $d_{t}\in\mathcal{D}_{\hat{i}}$ be the diagnosis that has been reached with $t$ iterations over $\mathcal{K}_{\hat{i}}$, where $t$ corresponds to the depth of node $d_{t}$ and so is less than or equal to the depth of $\mathcal{K}_{i}$. $d_{0}$ is the root node of $\mathcal{K}_{\hat{i}}$. For $d_{0}$, we apply the perception module to extract all observations in $R$ and explanation $\mathcal{E}_{0}$ to support $d_{0}$ as
$\hat{\mathcal{O}},\hat{\mathcal{E}_{0}}=W(d_{0},\mathcal{K}_{\hat{i}}).$ (3)
$\mathcal{K}_{\hat{i}}$ is supplied to facilitate the model to extract all observations for the following reasoning process. [^5] We used only pairs of an observation and a premise. We abuse $\mathcal{K}$ to mean this for notation simplicity.
Diagnosis $d_{t}$ identifies the set $\{d_{n}\}_{n}$ of its children and so the set $\mathcal{P}_{\hat{i}}(\{d_{n}\}_{n})=\{p\in\mathcal{P}_{i}|(p,d_{n})\in\mathcal{S}_{i},d_{n}\in\{d_{n}\}\}$ of premises that support $d_{n}$. Therefore, our reasoning module $V$ iteratively and greedily identifies the next step’s diagnosis (i.e., $d_{t+1}$) from $\{d_{n}\}_{n}$, making a rationalization for each deduction. That is, $V$ verifies whether there exist $o$’s in $\hat{\mathcal{O}}$ that supports one $d_{n}$. If $d_{n}$ is fully supported, $d_{n}$ is identified as $d_{t+1}$ for the $(t+1)$-th iteration, i.e.,
$d_{t+1},\hat{\mathcal{E}}_{t+1}=V(\hat{\mathcal{O}},\{d_{n}\},\mathcal{P}_{\hat{i}}(\{d_{n}\}_{n})).$ (4)
Otherwise, the reasoning module fails. $V$ is repeated until $d_{t^{\prime}}$ in $\mathcal{D}^{\star}_{\hat{i}}$ is found or it fails. In our annotation, each observation contributes to deducing only one $d_{t}$. Therefore, if an observation in ^$\hat{\mathcal{E}}_{t+1}$ is included in the preceding sets of explanations ^$\hat{\mathcal{E}}_{0}$ to ^$\hat{\mathcal{E}}_{t}$, the corresponding explanation in the preceding sets is removed.
5 Experiments
5.1 Experimental Setup
We assess the reasoning capabilities of 7 recent LLMs from diverse families and model sizes, including 5 instruction-tuned models that are openly accessible: LLama3 8B and 70B [AI@Meta, 2024], Zephyr 7B [Tunstall et al., 2023], Mistral 7B [Jiang et al., 2023], and Mixtral 8$\times$7B [Jiang et al., 2023]. We have also obtained access to private versions of the GPT-3.5 turbo [OpenAI, 2023b] and GPT-4 turbo [OpenAI, 2023a]  [^6] These two models are housed on a HIPPA-compliant instance within Microsoft Azure AI Studio. No data is transferred to either Microsoft or OpenAI. This secure environment enables us to safely conduct experiments with the MIMIC-IV dataset, in compliance with the Data Use Agreement., which are high-performance closed-source models. Each LLM is utilized to implement our baseline’s narrowing-down, perception, and reasoning modules. The temperature is set to 0. For computing evaluation metrics, we use LLama3 8B with few-shot prompts to make correspondences between $\mathcal{O}$ and $\hat{\mathcal{O}}$ as well as to verify a match between predicted and ground-truth explanations (refer to the supplementary material for more details).
5.2 Results
Comparison among LLMs. Table 3 shows the performance of our baseline built on top of various LLMs. We first evaluate a variant of our task that takes graph $\mathcal{G}=\{\mathcal{G}_{i}\}$ consisting of only procedural flow as external knowledge instead of $\mathcal{K}$. Comparison between $\mathcal{G}$ and $\mathcal{K}$ demonstrates the importance of supplying premises with the model and LLMs’ capability to make use of extensive external knowledge that may be superficially different from statements in $R$. Subsequently, some models are evaluated with our task using $\mathcal{K}$. In addition to the metrics in Section 3.5, we also adopt the accuracy of disease category $\textit{Acc}^{\text{cat}}$, which gives 1 when $\hat{i}=i$, as our baseline’s performance depends on it.
Table 3: Diagnostic reasoning ability of different LLMs under the proposed baseline method.
Diagnosis Observation Explanation
[…]
With $\mathcal{G}$ Zephyr 7B 0.274 0.151 0.123±${}_{\pm\text{0.200}}$ 0.115±${}_{\pm\text{0.166}}$ 0.092±${}_{\pm\text{0.108}}$ 0.071±${}_{\pm\text{0.139}}$ 0.014±${}_{\pm\text{0.037}}$
Mistral 7B 0.507 0.306 0.211±${}_{\pm\text{0.190}}$ 0.317±${}_{\pm\text{0.253}}$ 0.173±${}_{\pm\text{0.157}}$ 0.230±${}_{\pm\text{0.312}}$ 0.062±${}_{\pm\text{0.088}}$
Mixtral 8$\times$7B 0.413 0.237 0.147±${}_{\pm\text{0.165}}$ 0.266±${}_{\pm\text{0.261}}$ 0.124±${}_{\pm\text{0.138}}$ 0.144±${}_{\pm\text{0.268}}$ 0.029±${}_{\pm\text{0.056}}$
LLama3 8B 0.576 0.321 0.253±${}_{\pm\text{0.156}}$ 0.437±${}_{\pm\text{0.207}}$ 0.219±${}_{\pm\text{0.137}}$ 0.232±${}_{\pm\text{0.316}}$ 0.071±${}_{\pm\text{0.093}}$
LLama3 70B 0.752 0.540 0.277±${}_{\pm\text{0.146}}$ 0.537±${}_{\pm\text{0.192}}$ 0.256±${}_{\pm\text{0.142}}$ 0.395±${}_{\pm\text{0.320}}$ 0.112±${}_{\pm\text{0.110}}$
GPT-3.5 turbo 0.679 0.455 0.389±${}_{\pm\text{0.212}}$ 0.351±${}_{\pm\text{0.192}}$ 0.275±${}_{\pm\text{0.167}}$ 0.331±${}_{\pm\text{0.366}}$ 0.103±${}_{\pm\text{0.127}}$
GPT-4 turbo 0.772 0.572 0.446±${}_{\pm\text{0.207}}$ 0.491±${}_{\pm\text{0.180}}$ 0.371±${}_{\pm\text{0.186}}$ 0.475±${}_{\pm\text{0.363}}$ 0.199±${}_{\pm\text{0.181}}$
With $\mathcal{K}$ LLama3 8B 0.576 0.344 0.235±${}_{\pm\text{0.162}}$ 0.394±${}_{\pm\text{0.227}}$ 0.199±${}_{\pm\text{0.142}}$ 0.327±${}_{\pm\text{0.375}}$ 0.087±${}_{\pm\text{0.114}}$
LLama3 70B 0.735 0.581 0.262±${}_{\pm\text{0.146}}$ 0.501±${}_{\pm\text{0.208}}$ 0.236±${}_{\pm\text{0.131}}$ 0.463±${}_{\pm\text{0.374}}$ 0.125±${}_{\pm\text{0.117}}$
GPT-3.5 turbo 0.652 0.413 0.347±${}_{\pm\text{0.241}}$ 0.279±${}_{\pm\text{0.203}}$ 0.232±${}_{\pm\text{0.184}}$ 0.374±${}_{\pm\text{0.408}}$ 0.121±${}_{\pm\text{0.152}}$
GPT-4 turbo 0.781 0.614 0.431±${}_{\pm\text{0.207}}$ 0.458±${}_{\pm\text{0.187}}$ 0.353±${}_{\pm\text{0.170}}$ 0.633±${}_{\pm\text{0.338}}$ 0.247±${}_{\pm\text{0.201}}$
Table 4: Evaluation of diagnostic reasoning ability of LLMs when no external knowledge is provided.
Observation Explanation
[…]
With $\mathcal{D}^{\star}$ LLama3 8B 0.070 0.154±${}_{\pm\text{0.139}}$ 0.330±${}_{\pm\text{0.244}}$ 0.135±${}_{\pm\text{0.122}}$ 0.020±${}_{\pm\text{0.100}}$ 0.004±${}_{\pm\text{0.016}}$
LLama3 70B 0.502 0.257±${}_{\pm\text{0.150}}$ 0.509±${}_{\pm\text{0.213}}$ 0.237±${}_{\pm\text{0.145}}$ 0.138±${}_{\pm\text{0.209}}$ 0.034±${}_{\pm\text{0.054}}$
GPT-3.5 turbo 0.223 0.164±${}_{\pm\text{0.242}}$ 0.149±${}_{\pm\text{0.212}}$ 0.116±${}_{\pm\text{0.174}}$ 0.091±${}_{\pm\text{0.231}}$ 0.025±${}_{\pm\text{0.065}}$
GPT-4 turbo 0.636 0.461±${}_{\pm\text{0.206}}$ 0.482±${}_{\pm\text{0.160}}$ 0.378±${}_{\pm\text{0.174}}$ 0.186±${}_{\pm\text{0.221}}$ 0.074±${}_{\pm\text{0.090}}$
No Knowledge LLama3 8B 0.023 0.137±${}_{\pm\text{0.159}}$ 0.258±${}_{\pm\text{0.274}}$ 0.119±${}_{\pm\text{0.141}}$ 0.018±${}_{\pm\text{0.083}}$ 0.006±${}_{\pm\text{0.026}}$
LLama3 70B 0.037 0.246 ±${}_{\pm\text{0.148}}$ 0.504±${}_{\pm\text{0.222}}$ 0.227±${}_{\pm\text{0.148}}$ 0.022±${}_{\pm\text{0.093}}$ 0.007±${}_{\pm\text{0.030}}$
GPT-3.5 turbo 0.059 0.161±${}_{\pm\text{0.238}}$ 0.148±${}_{\pm\text{0.215}}$ 0.113±${}_{\pm\text{0.171}}$ 0.036±${}_{\pm\text{0.131}}$ 0.011±${}_{\pm\text{0.039}}$
GPT-4 turbo 0.074 0.410±${}_{\pm\text{0.208}}$ 0.443±${}_{\pm\text{0.191}}$ 0.324±${}_{\pm\text{0.182}}$ 0.047±${}_{\pm\text{0.143}}$ 0.019±${}_{\pm\text{0.058}}$
With $\mathcal{G}$, we can see that GPT-4 achieves the best performance in most metrics, especially related to observations and explanations, surpassing LLama3 70B by a large margin. In terms of accuracy (in both category and diagnosis levels), LLama3 70B is comparable to GPT-4. LLama3 70B also has a higher $\textit{Obs}^{\text{rec}}$ but low $\textit{Obs}^{\text{pre}}$ and $\textit{Obs}^{\text{comp}}$, which means that this model tends to extract many observations. Models with high diagnostic accuracy are not necessarily excel in finding essential information in long text (i.e., observations) and generating reasons (i.e., explanations).
[…]
LLMs may undergo inherent challenges for evaluation when no external knowledge is supplied. They may have the knowledge to diagnose but cannot make consistent observations and explanations that our task expects through $\mathcal{K}$. To explore this, we evaluate two settings: (1) giving $D^{\star}$ and (2) no knowledge is supplied to a model (shown in Table 4). The prompts used for this setup are detailed in the supplementary material. We do not evaluate the accuracy of disease category prediction as it is basically the same as Table 3. We can clearly see that with $\mathcal{D}^{\star}$, GPT-4’s diagnostic and observational scores are comparable to those of the task with $\mathcal{K}$, though explanatory performance is much worse. Without any external knowledge, the diagnostic accuracy is also inferior. [^7] We understand this comparison is unfair, as the prompts differ. We intend to give a rough idea about the challenge without external knowledge. The deteriorated performance can be attributed to inconsistent wording of diagnosis names, which makes evaluation tough. High observational scores imply that observations in $R$ can be identified without relying on external knowledge. There can be some cues to spot them.
Figure 5: Performance of LLama3 70B, GPT-3.5, and GPT-4 under different medical domains. We use the task with $\mathcal{G}$.
[…]
Table 5: Consistency of automated evaluation metrics with human judgments.
Model Observation Rationalization
LLama3 8B 0.887 0.801
GPT-4 turbo 0.902 0.836
Reliability of automatic evaluation. We randomly pick out 100 samples from DiReCT and their prediction by GPT-4 over the task with $\mathcal{G}$ to assess the consistency of our automated metrics to evaluate the observational and explanatory performance in Section 3.3 to human judgments. Three physicians joined this experiment. For each prediction $\hat{o}\in\hat{\mathcal{O}}$, they are asked to find a similar observation in ground truth $\mathcal{O}$. For explanatory metrics, they verify if each prediction $\hat{z}\in\hat{\mathcal{E}}$ for $\hat{o}\in\hat{\mathcal{O}}$ align with ground-truth $z\in\mathcal{E}$ corresponding to $o$. A prediction and a ground truth are deemed aligned for both assessments if at least two specialists agree. We compare LLama3’s and GPT-4’s judgments to explore if there is a gap between these LLMs. As summarized in Table 5, GPT-4 achieves the best results, with LLama3 8B also displaying a similar performance. From these results, we argue that our automated evaluation metrics are consistent with human judgments, and LLama3 is sufficient for this evaluation, allowing the cost-efficient option.
A prediction example. Figure 6 shows a sample generated by GPT-4. The ground-truth PDD of the input clinical note is Hemorrhagic Stroke. In this figure, purple, orange, and red indicate explanations only in the ground truth, only in prediction, and common in both, respectively; therefore, red is a successful prediction of an explanation, while purple and orange are a false negative and false positive. GPT-4 treats the observation of aurosis fugax as the criteria for diagnosing Ischemic Stroke. However, this observation only supports Suspected Stroke. Conversely, observation thalamic hematoma, which is the key indicator of Hemorrhagic Stroke, is regarded as a less important clue. Such observation-diagnosis correspondence errors lead to the model’s misdiagnosis. More samples are available in the supplementary material.
Figure 6: An example prediction for a clinical note with PDD of Hemorrhagic Stroke by GPT-4.
6 Conclusion and Limitations
We proposed DiReCT as the first benchmark for evaluating the diagnostic reasoning ability of LLMs with interpretability by supplying external knowledge as a graph. Our evaluations reveal a notable disparity between current leading-edge LLMs and human experts, underscoring the urgent need for AI models that can perform reliable and interpretable reasoning in clinical environments. DiReCT can be easily extended to more challenging settings by removing the knowledge graph from the input, facilitating evaluations of future LLMs.
Limitations. DiReCT encompasses only a subset of disease categories and considers only one PDD, omitting the inter-diagnostic relationships due to their complexity—a significant challenge even for human doctors. Additionally, our baseline may not use optimal prompts, chain-of-thought reasoning, or address issues related to hallucinations in task responses. Our dataset is solely intended for model evaluation but not for use in clinical environments. The use of the diagnostic knowledge graph is also limited to serving merely as a part of the input. Future work will focus on constructing a more comprehensive disease dataset and developing an extensive diagnostic knowledge graph.
Acknowledgments and Disclosure of Funding
This work was supported by World Premier International Research Center Initiative (WPI), MEXT, Japan. This work is also supported by JSPS KAKENHI 24K20795 and Dalian Haichuang Project for Advanced Talents.
Appendix A Details of DiReCT
A.1 Data Statistics
Table 6: Disease statistics of DiReCT.
Domains Categories # samples $|\mathcal{D}_{i}|$ $|\mathcal{D}^{\star}_{i}|$ References
Cardiology Acute Coronary Syndromes 65 6 3 [Byrne et al., 2024]
Aortic Dissection 14 3 2 [Members et al., 2022]
Atrial Fibrillation 10 3 2 [Joglar et al., 2024]
Cardiomyopathy 9 5 4 [Ommen et al., 2020]
Heart Failure 52 6 3 [Heidenreich et al., 2022]
Hyperlipidemia 2 2 1 [Su et al., 2021]
Hypertension 32 2 1 [Unger et al., 2020]
Gastroenterology Gastritis 27 5 3 [Shah et al., 2021]
Gastroesophageal Reflux Disease 41 2 1 [Gyawali et al., 2024]
Peptic Ulcer Disease 28 3 2 [Kavitt et al., 2019]
Upper Gastrointestinal Bleeding 7 2 1 [Barkun et al., 2019]
Neurology Alzheimer 10 2 1 [McKhann et al., 1984]
Epilepsy 8 3 2 [Igaku-Shoin-Ltd., 2018]
Migraine 4 3 2 [Lipton et al., 2001]
Multiple Sclerosis 27 6 4 [Lublin, 2005]
Stroke 28 3 2 [Kleindorfer et al., 2021]
Pulmonology Asthma 13 7 5 [Qaseem et al., 2011]
COPD 19 6 4 [Gupta et al., 2013]
Pneumonia 20 4 2 [Olson and Davis, 2020]
Pulmonary Embolism 35 5 3 [Konstantinides et al., 2020]
Tuberculosis 5 3 2 [Lewinsohn et al., 2017]
Endocrinology Adrenal Insufficiency 20 4 3 [Charmandari et al., 2014]
Diabetes 13 4 2 [ElSayed et al., 2023]
Pituitary 12 4 3 [Tritos and Miller, 2023]
Thyroid Disease 10 6 4 [AlexanderErik et al., 2017]
Table 6 provides a detailed breakdown of the disease categories included in DiReCT. The column labeled # samples indicates the number of data points. The symbols $|\mathcal{D}_{i}|$ and $|\mathcal{D}^{\star}_{i}|$ denote the total number of diagnoses (diseases) and PDDs, respectively. Existing guidelines for diagnosing diseases were used as References, forming the foundation for constructing the diagnostic knowledge graphs. As some premise may not included in the referred guidelines. During annotation, physicians will incorporate their own knowledge to complete the knowledge graph.
A.2 Structure of Knowledge Graph
The entire knowledge graph, denoted as $\mathcal{K}$, is stored in separate JSON files, each corresponding to a specific disease category $i$ as $\mathcal{K}_{i}$. Each $\mathcal{K}_{i}$ comprises a procedural graph $\mathcal{G}_{i}$ and the corresponding premise $p$ for each disease. As illustrated in Figure 7, the procedural graph $\mathcal{G}_{i}$ is stored under the key "Diagnostic" in a dictionary structure. A key with an empty list as its value indicates a leaf diagnostic node as $d^{\star}$. The premise for each disease is saved under the key of "Knowledge" with the corresponding disease name as an index. For all the root nodes (e.g., Suspected Heart Failure), we further divide the premise into "Risk Factors", "Symptoms", and "Signs". Note that each premise is separated by ";".
Figure 7: A sample of knowledge graph for Heart Failure. Each premise under the key of "Knowledge" is separated with ";".
A.3 Annotation and Tools
We have developed proprietary software for annotation purposes. As depicted in Figure 8, annotators are presented with the original text as observations $o$ and are required to provide rationales ($z$) to explain why a particular observation $o$ supports a disease $d$. The left section of the figure, labeled Input1 to Input6, corresponds to different parts of the clinical note, specifically the chief complaint, history of present illness, past medical history, family history, physical exam, and pertinent results, respectively. Annotators will add the raw text into the first layer by left-clicking and dragging to select the original text, then right-clicking to add it. After each observation, a white box will be used to record the rationales. Finally, a connection will be made from each rationale to a disease, represented in a grey box. The annotation process strictly follow the knowledge graph. Both the final annotation and the raw clinical note will be saved in a JSON file. We provide the code to compile these annotations and detailed instructions for using our tool on GitHub.
Figure 8: Demonstration of our annotation tool.
A.4 Access to DiReCT
Implementation code and annotation tool are available through https://github.com/wbw520/DiReCT. Data will be released through PhysioNet due to safety issues according to the license of MIMIC-IV (PhysioNet Credentialed Health Data License 1.5.0). We will use the same license for DiReCT. The download link will be accessible via GitHub. We confirm that this GitHub link and data link are always accessible. We confirm that we will bear all responsibility in case of violation of rights.
Appendix B Implementation of Baseline Method
B.1 Prompt Settings
Table 7: Prompt for narrowing-down module.
Input Prompt
Suppose you are one of the greatest AI scientists and medical expert. Let us think step by step.
You will review a clinical ’Note’ and your ’Response’ is to diagnose the disease that the patient have for this admission.
All possible disease options are in a list structure: {disease_option}.
Note that you can only choose one disease from the disease options and directly output the origin name of that disease.
Now, start to complete your task.
Don’t output any information other than your ’Response’.
’Note’:
{note}
Your ’Response’:
Table 8: Prompt for perception module.
Input Prompt
Suppose you are one of the greatest AI scientists and medical expert. Let us think step by step.
You will review a part of clinical "Note" from a patient.
The disease for which the patient was admitted to hospital this time is {disease}.
Your task is to extract the original text as confidence "Observations" that lead to {disease}.
Here are some premise for the diagnosis of this disease category. You can refer them for your task. Premise are: {premise}
Note that you also need to briefly provide the "Reason" for your extraction.
Note that both "Observations" and "Reason" should be string.
Note that your "Response" should be a list structure as following
: [["Observation", "Reason"], ……, ["Observation", "Reason"]]
Note that if you can’t find any "Observation" your "Response" should be: [].
Now, start to complete your task.
Note that you should not output any information other than your "Response".
"Note":
{note}
Note that you should not output any information other than your "Response".
Your "Response":
Table 9: Prompt for reasoning module.
Input Prompt
Suppose you are one of the greatest AI scientists and medical expert. Let us think step by step.
You will receive a list of "Observations" from a clinical "Note". These "Observations" are possible support to diagnose {disease}.
Based on these "Observations", you need to diagnose the "Disease" from the following options: {disease_option}.
Here are some golden standards to discriminate diseases. You can refer them for your task. Golden standards are: {premise}
Note that you can only choose one "Disease" from the disease options and directly output the name in disease options.
Note that you also required to select the "Observations" that satisfy the golden standard to diagnose the "Disease" you choose.
Note that you also required to provide the "Reason" for your choice.
Note that your "Response" should be a list structure as following
:[["Observation", "Reason", "Disease"], ……, ["Observation", "Reason", "Disease"]]
Note that if you can’t find any "Observation" to support a disease option, your "Response" should be: None
Now, start to complete your task.
Note that you should not output any information other than your "Response".
"Observations":
{observation}
Note that you should not output any information other than your "Response".
Your "Response":
In this section, we demonstrate the prompt we used for each module (From Table 7-9 for narrowing-down, perception, and reasoning module, respectively).
In Table 7, {disease_option} is the name for all disease categories, and {note} is the content for the whole clinical note. The response for the model is the name of a possible disease $\hat{i}$.
In Table 8, {disease} is the disease category name predicted in narrowing-down. The content marked blue is the premise, which is only provided during the $\mathcal{k}$ setting. In this module, {premise} is offered with all information in the knowledge graph. Different to narrowing-down, {note} is implemented for each clinical data $R=\{r\}$ and the outputs are combined together for $\hat{\mathcal{O}}$ and $\hat{\mathcal{E}}$.
In Table 9, {disease} is the disease category name and {disease_option} is consisted by the children nodes $\{d_{n}\}_{n}$. Similarly, the premise on the blue is only available for the $\mathcal{k}$ setting. It provides the premise that are criteria for the diagnosis of each children node. {observation} is the extracted $\hat{\mathcal{O}}$ in previous step. We provide all the prompts and the complete implementation code on GitHub.
B.2 Details of Automatic Evaluation
The automatic evaluation is realized by LLama3 8B. We demonstrate the prompt for this implement in Table 10 (for observation) and Table 11 (for rationalization). Note that we do not use few-shot samples for the evaluation of observation. In Table 10, {gt_observation} and {pred_observation} are from model prediction and ground-truth. As this is a simple similarity comparison task to discriminate whether the model finds similar observations to humans, LLama3 itself have such ability. We do not strict to exactly match due to the difference in length of extracted raw text (as long as the observation expresses the same description). In Table 11, {gt_reasoning} and {pred_reasoning} are from model prediction and ground-truth. We require the rationale to be complete (content of the expression can be understood from the rationale alone) and meaningful; therefore, we provide five samples for this evaluation. We also provide all the prompts and the complete implementation code on GitHub.
Table 10: Prompt for evaluation of observation.
Input Prompt
Suppose you are one of the greatest AI scientists and medical expert. Let us think step by step.
You will receive two "Observations" extracted from a patient’s clinical note.
Your task is to discriminate whether they textually description is similar?
Note that "Response" should be one selection from "Yes" or "No".
Now, start to complete your task.
Don’t output any information other than your "Response".
"Observation 1": {gt_observation}
"Observation 2": {pred_observation}
Your "Response":
Table 11: Prompt for evaluation of rationalization.
Input Prompt
Suppose you are one of the greatest AI scientists and medical expert. Let us think step by step.
You will receive two "Reasoning" for the explanation of why an observation cause a disease.
Your task is to discriminate whether they explain a similar medical diagnosis premise?
Note that "Response" should be one selection from "Yes" or "No".
Here are some samples:
Sample 1:
"Reasoning 1": Facial sagging is a classic symptom of stroke
"Reasoning 2": Indicates possible facial nerve palsy, a common symptom of stroke
"Response": Yes
Sample 2:
"Reasoning 1": Family history of Diabetes is an important factor
"Reasoning 2": Patient’s mother had a history of Diabetes, indicating a possible genetic predisposition to stroke
"Response": Yes
Sample 3:
"Reasoning 1": headache is one of the common symptoms of HTN
"Reasoning 2": Possible symptom of HTN
"Response": No
Sample 4:
"Reasoning 1": Acute bleeding is one of the typical symptoms of hemorrhagic stroke
"Reasoning 2": The presence of high-density areas on Non-contrast CT Scan is a golden standard for Hemorrhagic Stroke
"Response": No
Sample 5:
"Reasoning 1": Loss of strength on one side of the body, especially when compared to the other side, is a common sign of stroke
"Reasoning 2": Supports ischemic stroke diagnosis
"Response": No
Now, start to complete your task.
Don’t output any information other than your "Response".
"Reasoning 1": {gt_reasoning}
"Reasoning 2": {pred_reasoning}
Your "Response":
B.3 Prediction Samples
Figure 9 and 10 shows two sample generated by GPT-4. The ground-truth PDD of the input clinical note is Gastroesophageal Reflux Disease (GERD) and Heart Failure (HF). In these figure, purple, orange, and red indicate explanations only in the ground truth, only in prediction, and common in both, respectively; therefore, red is a successful prediction of an explanation, while purple and orange are a false negative and false positive.
Figure 9: An example prediction for a clinical note with PDD of GERD by GPT-4
Figure 10: An example prediction for a clinical note with PDD of HF by GPT-4
In Figure 9, we can observe that GPT-4 can find the key observation for the diagnosis of GERD, which is consistent with human in both observation and rationale. However, it still lacks the ability to identify all observations and establish accurate relationships for diseases. In Figure 10, the model’s predictions do not align well with those of a human doctor. Key observations, such as the relationships between BNP and LVEF, are incorrectly identified, leading to a final misdiagnosis.
B.4 Experiments for No Extra Knowledge
Table 12: Prompt for $\mathcal{D}^{\star}$ setting.
Input Prompt
Suppose you are one of the greatest AI scientists and medical expert. Let us think step by step.
You will review a clinical ’Note’ and your ’Response’ is to diagnose the disease that the patient have for this admission.
All possible disease options are in a list structure: {disease_options}.
Note that you can only choose one disease from the disease options and directly output the origin name of that disease.
Now, start to complete your task.
Don’t output any information other than your ’Response’.
’Note’:
{note}
Your ’Response’:
Table 13: Prompt for no knowledge setting.
Input Prompt
Suppose you are one of the greatest AI scientists and medical expert. Let us think step by step.
You will review a clinical ’Note’ and your ’Response’ is to diagnose the disease that the patient have for this admission.
Note that you can only give one disease name and directly output the name of that "Disease".
Now, start to complete your task.
Don’t output any information other than your ’Response’.
’Note’:
{note}
Your ’Response’:
We demonstrate the prompt used for $\mathcal{D}^{\star}$ and no knowledge settings in Table 12 and Table 13, respectively. {note} is the text of whole clinical note and {note} in Table 12 is the name of all leaf node $\mathcal{D}^{\star}$.
B.5 Experimental Settings
All experiments are implemented with a temperature value of 0. All close sourced models are implemented in a local server with 4 NVIDIA A100 GPU.
Appendix C Failed Attempts on DiReCT
In this section, we discuss some unsuccessful attempts during the experiments.
Extract observation from the whole clinical note. We try to diagnose the disease and extract observation, and the corresponding rationale using the prompt shown in Table 14. The {note} is offered by the whole content in the clinical note. We find that even though the model can make the correct diagnosis, only a few observations can be extracted (no more than 4), which decreases the completeness and faithfulness.
Table 14: Prompt for extracting observation in one step.
Input Prompt
Suppose you are one of the greatest AI scientists and medical expert. Let us think step by step.
You will review a clinical ’Note’, and your ’Response’ is to diagnose the disease that the patient has for this admission.
All possible disease options are in a list structure: {disease_options}.
Note that you can only choose one disease from the disease options and directly output the origin name of that disease.
Note that you also need to extract original text as confidence "Observations" that lead to the "Disease" you selected.
Note that you should extract all necessary "Observation".
Note that you also need to briefly provide the "Reason" for your extraction.
Note that both "Observations" and "Reason" should be string.
Note that your "Response" should be a list structure as following
:[["Observation", "Reason", "Disease"], ……, ["Observation", "Reason", "Disease"]]
Now, start to complete your task.
Don’t output any information other than your ’Response’.
’Note’
:{note}
Your ’Response’:
End-to-End prediction. We also try to output the whole reasoning process in one step (without iteration) when given observations. We show our prompt in Table 15. We find that using such a prompt model can not correctly recognize the relation between observation, rationale, and diagnosis.
Table 15: Prompt for extracting observation in one step.
Input Prompt
Suppose you are one of the greatest AI scientists and medical expert. Let us think step by step.
You will receive a list of "Observations" from a clinical "Note" for the diagnosis of stroke.
Here is the diagnostic route of stroke in a tree structure:
-Suspected Stroke
     -Hemorrhagic Stroke
     -Ischemic Stroke
Here are some premise for the diagnosis of this disease. You can refer them for your task. Premise are: {premise}
Based on these "Observations", starting from the root disease, your target is to diagnose one of the leaf disease.
Note that you also required to provide the "Reason" for your reasoning.
Note that your "Response" should be a list structure as following
:[["Observation", "Reason", "Disease"], ……, ["Observation", "Reason", "Disease"]]
Note that if you can’t find any "Observation" to support a disease option, your "Response" should be: None
Now, start to complete your task.
Note that you should not output any information other than your "Response".
"Observations":
{observation}
Note that you should not output any information other than your "Response".
Your "Response":
Appendix D Ethical Considerations
Utilizing real-world EHRs, even in a de-identified form, poses inherent risks to patient privacy. Therefore, it is essential to implement rigorous data protection and privacy measures to safeguard sensitive information, in accordance with regulations such as HIPAA. We strictly adhere to the Data Use Agreement of the MIMIC dataset, ensuring that the data is not shared with any third parties. All experiments are implement on a private server. The access to GPT is also a private version.
AI models are susceptible to replicating and even intensifying the biases inherent in their training data. These biases, if not addressed, can have profound implications, particularly in sensitive domains such as healthcare. Unconscious biases in healthcare systems can result in significant disparities in the quality of care and health outcomes among different demographic groups. Therefore, it is imperative to rigorously examine AI models for potential biases and implement robust mechanisms for ongoing monitoring and evaluation. This involves analyzing the model’s performance across various demographic groups, identifying any disparities, and making necessary adjustments to ensure equitable treatment for all. Continual vigilance and proactive measures are essential to mitigate the risk of biased decision-making and to uphold the principles of fairness and justice in AI-driven healthcare solutions.